OpenStudy (alyssajobug):

Identify the axis of symmetry for the graph of g(x) = 5(x - 2)^2 + 4

OpenStudy (alyssajobug):

I tried it multiple times, at first I got -2, then I thought about it more and got 20, then finally I switched to using PEMDAS and got 100 and none were right so far... HELP?!?!

OpenStudy (amoodarya):

y=a(x-b)^2 +c x=b is the axis of symmetry

OpenStudy (amoodarya):

like this y=2(x-3)^2 +11 x=3 is it (axis of symmetry) now can you go on ?

OpenStudy (alyssajobug):

so no negitive sign goes into the axis of symmetry?

OpenStudy (alyssajobug):

OH! IT WORKED WITH POSITIVE 2!!!!!

OpenStudy (amoodarya):

if you write the equation in standard form like that y=a(x-b)^2 +c the zero of (x-b)^2 is axis of symmetry so (x-2)^2=0 then (x-2)=0 then x=2

OpenStudy (alyssajobug):

So, in: y=a(x-b)^2 +c x=b is the axis of symmetry equasion, can you break down what each mean? My lessons are a joke and yet they expect me to learn all this!

OpenStudy (alyssajobug):

How do I graph like that? That could come in handy!!!

OpenStudy (amoodarya):